OpenStudy (anonymous):

If a triangle has as an area of 144 ft2 and a height of 24 ft, then what is the base?

OpenStudy (anonymous):

formula for area of triangle= 1/2 * base*height so, here, 144= 1/2 * 24* height so, height= 144/ (1/2 * 24)=12

jhonyy9 (jhonyy9):

144=b*h/2 144=b*24/2 144=b*12 b=12